OpenStudy (anonymous):

Futures Question: Suppose that the risk-free rate (continuously compounded) is 5% p.a. The stock-market index is currently at £ 5,000 and the futures price for a contract deliverable in 3 months is £ 5,050. Question - what arbitrage opportunities does this create? how would you exploit them?

OpenStudy (anonymous):

Firstly we re-calculate the interest rate: 0,05/12*3= 0,0125 Then we compound the sum: 5000*(1+0,0125)=5062 The arbitrage opportunity is 5062-5050=12 We can exploit AO by selling above mentioned futures contract and investing into the stock-market index for 3 months.

OpenStudy (anonymous):

Hmm, 2 mistakes there (one big, one small), anyone else?

OpenStudy (anonymous):

I see .. Its continuously compounded not annual percentage rate. Then its ... 3 month interest rate: [(\sqrt[12]{1,05})^{3}-1=0,0123\] Then we compound the sum: 5000*(1+0,0123)=5062 The arbitrage opportunity is 5062-5050=12 We discount the profit back to t=0: 12*\[e^{-0,0123}\]=11,85

OpenStudy (anonymous):

correct on the continous compounding :-) Think of the pair of trades you suggested for the AO though ...

OpenStudy (anonymous):

Its vice-versa. Futures contract allows you to buy stock index for less money :)

OpenStudy (anonymous):

Good answer. - Buy the stock index future (it's underpriced at £5,050) - Short the index spot in the cash market - Invest the proceeds in Treasuries (£5,000) for 3 months. At maturity, at 3months: - Cash out the risk-free investment at £5,062 - take delivery of the index, for £5,050; close out the short index position (now you're flat). - pocket the $11.85 difference. :-) risk-free Arbitrage opportunity.